# Fullness of $q$-Araki-Woods Factors

### v1

@@ 3. Technical analysis
This section contains the rigorous technical analysis, where we will prove a series of Lemmas and Propositions about various convergence results. The most crucial results are Propositions 3.9, 3.4 and 3.12. We recommend that readers first review the statements of these Propositions and then proceed directly to the proof of the main theorem in the next section to understand the reasons behind these calculations. The notations used in Proposition 3.9 are mentioned just before Lemma 3.7.
@@
where $X$ is an operator bounded by $\frac{C_{q}^{6}}{1-|q|}(1-q)^{-n-\frac{\ell}{2}}\|\xi\|_{U}^{2n}\|\eta\|_{U}^{\ell}$. By multiplying both side from left by $c^{*}(\xi)^{m-n}$ and using the $q$-commutation relation $c^{*}(\xi)^{m-n}c(\eta)^{\ell}=q^{(m-n)\ell}c(\eta)^{\ell}c^{*}(\xi)^{(m-n)}$ since $\langle\xi,\eta\rangle_{U}=0$, we get
@@
It is very important that we multiply the operator $T_{k}$ from the right in Proposition 3.4 and not from the left. In fact, the operator $\tilde{S}_{n}=\frac{1}{n}\sum_{k=0}^{n}(1-q)^{k}\tilde{T}_{k}c(\xi)^{k}c(\eta)^{k}$ need not converge to 00 for fixed unit vectors $\xi,\eta\in\mathsf{H}_{U}$ with $\langle\xi,\eta\rangle_{U}=0$ and bounded operators $\tilde{T}_{k}$. See for example Proposition 3.9 below. However, for special operators $\tilde{T}_{k}$, we show in Proposition 3.4 that $\tilde{S}_{n}$ converges to 00.
Lemma 3.6.
For any non-zero vector $\xi\in\mathsf{H}_{U}$, the sequence
$Z_{n}=(1-q)^{n}\|\xi\|_{U}^{-2n}c^{*}(\xi)^{n}c(\xi)^{n}$
is convergent in norm to some operator $Z$. Moreover $Z$ is positive and invertible such that $d_{\infty}\leq Z\leq 1$ for $q\geq 0$, and $\frac{d_{\infty}}{1-q}\leq Z\leq 1-q$ for $q\leq 0$.
Proof.
The proof follows the same lines of [BMRW, Lemma 3.4]. ∎
We now restrict our attention to a two-dimensional invariant subspace of a given representation $\{U_{t}\}_{t\in\mathbb{R}}$. Let $\mathsf{K}_{\mathbb{R}}=\mathbb{R}^{2}$ be an invariant subspace of $U_{t}$ such that ${U_{t}}_{|_{\mathsf{K}_{\mathbb{R}}}}$ is ergodic i.e. has no fixed nonzero vector. Then, as discussed in Section 2, there is a scalar $\lambda\in(0,1)$ such that
@@
Moreover, the (right) conjugate of $e$ and $\bar{e}$ with respect to the decomposition in $\mathsf{H}_{\mathbb{R}}^{\prime}+i\mathsf{H}_{\mathbb{R}}^{\prime}$ is given as follow:
@@
We briefly sketch the proof from [BMRW, Lemma 3.8] in order to set some notations and have some useful observations. Note that both the sequences have the same norm (by using the antilinear involution $J_{\varphi}$). We consider the first sequence which by Wick formula is given by
@@
where $q_{k,\ell}^{(n)}=\sum_{J_{1},J_{2}}q^{c(J,J^{c})}$ with the sum running over all possible subsets $J_{1}\subseteq\{1,\ldots,n\}$ of cardinality $k$ and $J_{2}\subseteq\{n+1,\ldots,2n\}$ of cardinality $\ell$ such that $J=J_{1}\cup J_{2}$. Then we have $|q_{k,\ell}^{(n)}|\leq C_{q}^{2}|q|^{(n-k)\ell}$ and
@@
(6) $s_{n}$ $=\frac{1}{n}\sum_{m=1}^{n}\sum_{0\leq k<\ell\leq m}q^{(m-k)\ell}\lambda^{\frac{k-\ell}{2}}+\frac{1}{n}\sum_{m=1}^{n}\sum_{1\leq\ell\leq k\leq m-1}q^{(m-k)\ell}\lambda^{\frac{k-\ell}{2}}$
@@
where $C(q,\lambda,n_{0})$ is a constant independent from $n$, and where we have also used the fact that the sequence $\left\{(m+1)\frac{q^{m}}{2q^{n_{0}}-q^{m}}\right\}_{m\geq 1}$ converges to 00 as $m\to\infty$ and hence its Cesaro sum also converges to 00. On the other hand, the second quantity in (7) is estimated as follows:
@@
as the sequence $\{\frac{q^{k}}{1-q^{k}}\}_{k\geq 1}$ converges to 00 as $k\to\infty$, and so does its Cesaro sum. Thus we have proved that $s_{n}\to 0$ as $n\to\infty$. ∎
@@
First check that the series defined in (8) above is norm convergent. Indeed by using Lemma 2.5 we have
$\|S\|$ $\leq\sum_{k=0}^{\infty}\;\frac{d_{\infty}}{d_{k}}(1-q)^{k}\;\lambda^{\frac{k}{2}}\|c(\bar{e})^{k}\|\;\|T\|\;\|c(e)^{k}\|$
@@
where $|q^{(m)}_{k,\ell}|\leq C_{q}^{2}|q|^{(m-k)\ell}$. Let us understand the terms $c^{*}(\bar{e})^{m-\ell}c(\bar{e})^{m}c(e)^{m}$ in the sum above. Use Corollary 3.2 to write
@@
$\leq\frac{C_{q}^{9}A_{q}}{n^{2}}\sum_{n\geq m>m^{\prime}\geq 1}(1-q)^{2m+2m^{\prime}}\sum_{\ell=0}^{m}\sum_{\ell^{\prime}=0}^{m^{\prime}}|q|^{\ell^{\prime}}\frac{d_{m}d_{m^{\prime}}(1-q)^{-m-m^{\prime}}}{d_{\ell}d_{\ell^{\prime}}(1-q)^{-\ell-\ell^{\prime}}}\lambda^{\frac{m+m^{\prime}-\ell-\ell^{\prime}}{2}}(1-q)^{-m-m^{\prime}-\ell-\ell^{\prime}}\quad\quad\quad(\text{as }q_{m,\ell}^{(m)}\leq C_{q}^{2})$
@@
where $T_{p}=(1-q)^{p}\lambda^{\frac{p}{2}}c^{*}(e)^{p}c(e)^{p}$ for any $p\geq 1$. Let $T=\lim_{p\to\infty}T_{p}$ in norm, by using Lemma 3.6. Also for $0\leq k\leq m$, it is straightforward to verify that $q_{k,0}^{(m)}=\begin{pmatrix}m\\ k\end{pmatrix}_{q}$ (see for example, [BKS, Corollary 2.8]); so that for each $k\geq 0$, we have
@@
In particular, the Cesaro sum of the above sequence has the same limit; hence we get
@@
where we have used the fact that $\|T^{-1}\|\;\|T\|\leq\frac{1}{d_{\infty}}\leq C_{q}$ from Lemma 3.6. In particular, $d_{\infty}^{-1}T^{-1}S$ is invertible if $C_{q}^{4}\frac{\lambda^{\frac{1}{2}}}{{1-\lambda^{\frac{1}{2}}}}<1$, which is satisfied by the given value of $\lambda$ in the hypothesis. ∎
@@
and since $T^{-1}\leq 2d_{\infty}^{-1}\leq 2C_{q}$ (Lemma 3.6), we must have
@@
2. Although we have shown that the Cesaro sum of the sequence $\{W(\bar{e}^{m}e^{m})c(\bar{e})^{m}c(e)^{m}\}_{m\geq 1}$ converges, the sequence itself does not converge. For example, this can be checked for $q=0$ where the main obstruction is that the sequence $\{\sum_{\ell=0}^{m}c(\bar{e})^{m}c(e)^{\ell}c^{*}(\bar{e})^{m-\ell}c(\bar{e})^{m}c(e)^{m}\}_{m\geq 1}$ is not convergent. This is one very important instance for us to consider Cesaro sum of such bounded sequence.
We now prove some convergence results involving both the left and right creation/annihilation operators. For the purpose, we introduce the operator $Q$ defined by
@@
We may and we will assume that $\langle\xi,\eta\rangle_{U}\neq 0$, otherwise there is nothing to prove since $r^{*}(\xi)c(\eta)=c(\eta)r^{*}(\xi)$ whenever $\langle\xi,\eta\rangle_{U}=0$. By replacing $\xi$ with $\frac{\xi}{\langle\xi,\eta\rangle_{U}}$ (or $\eta$ with $\frac{\eta}{\langle\xi,\eta\rangle_{U}})$, we will further assume that $\langle\xi,\eta\rangle_{U}=1$.
@@
(12) $=q^{n-1}c(\eta)^{n-1}Q+c(\eta)\left(c(\eta)^{n-1}r^{*}(\xi)+[n-1]_{q}c(\eta)^{n-2}Q\right)$
(13) $=c(\eta)^{n}r^{*}(\xi)+[n]_{q}c(\eta)^{n-1}Q$
as required. Now we prove the result for general $m$ and $n$ such that $n\geq m$. The case $n<m$ follows similarly, which we will not expand. So now we fix $n$. The proof proceeds by induction on $m$. For $m=1$, the expression is proved above. By assuming the expression to be true for $m-1$, we calculate
@@
The following proposition can be considered a stronger version of [BMRW, Lemma 3.11] in the sense of norm convergence result over weak convergence.
@@
and $P_{\Omega}$ denotes the projection onto the vacuum vector $\Omega$.
@@
where the sum is divided over $(k,\ell=m)$, $(k,\ell=0)$ and the rest i.e. $X_{n}$ is the operator
@@
Thus we have divided $R_{n}$ as a sum of three terms, where we calculate norm limit of each term. To get the limit of the first term, first verify that $Q^{2m}$ converges to $P_{\Omega}$ as $m\to\infty$, and therefore the average $\frac{1}{n}\sum_{m=1}^{n}d_{m}^{2}Q^{2m}$ converges to $d_{\infty}^{2}P_{\Omega}$ in norm. The second term
@@
The sequence $\{m^{2}q^{m}\}_{m\geq 1}$ converges to 00 as $m\to\infty$, so does its average; hence we conclude that $X_{n}$ converges to 00 in norm as $n\to\infty$. ∎
@@
where $q_{k,\ell}^{(m)}$ is the same scalars as in the proof of Lemma 3.7.
@@ 4. Main result
[…]
Theorem 4.1.
The $q$-Araki-Woods factor $\mathsf{M}_{q}:=\Gamma_{q}(\mathsf{H}_{\mathbb{R}},U_{t})$ is full as soon as $\dim\mathsf{H}_{\mathbb{R}}\geq 2$.
Proof.
If $\mathsf{H}_{\mathbb{R}}$ is finite-dimensional, then $\mathsf{M}_{q}$ is full by [KSW, Theorem 2.7]. If $\mathsf{H}_{\mathbb{R}}$ has a non-zero weakly mixing part or if the point spectrum $\sigma_{p}(A)$ of eigenvalues of the analytic generator $A$ of $U_{t}$ is infinite and bounded, then $\mathsf{M}_{q}$ is full by [HI, Theorem 6.2]. So assume that $\mathsf{H}_{\mathbb{R}}$ is infinite-dimensional such that there is a strictly decreasing sequences of the set $\sigma_{p}(A)$ of eigenvalues of $A$ going to 00.
@@
Assume that $\lambda<(1+C_{q}^{4})^{-2}$ so that the norm limit $S$ of $S_{n}$ exists and is invertible by Proposition 3.9.
Now let $\omega$ be a fixed non-principal ultrafilter. Let $(X_{k})_{\omega}\in\mathsf{M}_{q}^{\prime}\cap\mathsf{M}_{q}^{\omega}$. By replacing each $X_{k}$ by $X_{k}-\varphi(X_{k})$, we may assume that $\varphi(X_{k})=0$ for all $k\geq 1$. In particular, we have $X_{k}\Omega,X_{k}^{*}\Omega\in\mathcal{F}_{q}(\mathsf{H}_{U})\ominus\mathbb{C}\Omega$ for all $k\geq 1$. We will show that $(X_{k})_{\omega}=0$ in $\mathsf{M}_{q}^{\omega}$ i.e. $\lim_{k\to\omega}\|X_{k}\Omega\|_{q}=0=\lim_{k\to\omega}\|X_{k}^{*}\Omega\|_{q}$. We observe
@@
where we could exchange the limit because $S_{n}S^{-1}$ converges in norm to identity as $n\to\infty$. Now we use the expressions $XX_{k}-X_{k}X\to 0$**-strongly as $k\to\omega$ for all $X\in\mathsf{M}_{q}$, and $X_{k}X^{\prime}-X^{\prime}X_{k}=0$ for all $X^{\prime}\in\mathsf{M}_{q}^{\prime}$ to get
@@
Propositions 3.4 and 3.12 tell us that the operators $R_{n}$ (or $R_{n}^{*}$) when restricted to $\mathcal{F}_{q}(\mathsf{H}_{U})\ominus\mathbb{C}\Omega$ converge to 00 in norm as $n\to\infty$. Since $X_{k}\Omega\in\mathcal{F}_{q}(\mathsf{H}_{U})\ominus\mathbb{C}\Omega$, it follows that
@@
where $L=\sup_{k\geq 1}\|X_{k}\Omega\|^{2}$, and we have used the upper bound for $\|S^{-1}\|$ from Remark 3.10. By our assumption, we have a sequence of such $\lambda$’s which converges to 00; hence it follows that $\lim_{k\to\omega}\|X_{k}\Omega\|_{q}=0$. A similar argument shows that $\lim_{k\to\omega}\|X_{k}^{*}\Omega\|_{q}=0$. This completes the proof. ∎
Acknowledgments. We thank Zhiyuan Yang (Texas A&M University) for going through the initial draft, and providing valuable suggestions; particularly, he suggested a shortened proof of Proposition 3.12 by utilizing the operator $Q$. This work was partially completed when M.K. was visiting S.W. at Harbin Institute of Technology in Aug-Sep 2023 for a month. M.K. extends appreciation to HIT for their warm hospitality and to S.W. for the invitation and financial support. M.K. was partially supported by the National Science Center (NCN) grant no. 2020/39/I/ST1/01566, as well as by NCN grant no. 2021/43/D/ST1/01446. S.W. was partially supported by the NSF of China (No.12031004, No.12301161) and the Fundamental Research Funds for the Central Universities.

### v2

@@ 3. Technical analysis
This section contains the rigorous technical analysis, where we will prove a series of Lemmas and Propositions about various convergence results. The most crucial results are Propositions 3.8, 3.4 and 3.11. We recommend that readers first review the statements of these Propositions and then proceed directly to the proof of the main theorem in the next section to understand the reasons behind these calculations. The notations used in Proposition 3.8 are mentioned just before Lemma 3.6.
@@
where $X$ is an operator bounded by $\frac{C_{q}^{6}}{1-|q|}(1-q)^{-n-\frac{\ell}{2}}\|\xi\|_{U}^{2n}\|\eta\|_{U}^{\ell}$. By multiplying both sides from the left by $c^{*}(\xi)^{m-n}$ and using the $q$-commutation relation $c^{*}(\xi)^{m-n}c(\eta)^{\ell}=q^{(m-n)\ell}c(\eta)^{\ell}c^{*}(\xi)^{(m-n)}$ since $\langle\xi,\eta\rangle_{U}=0$, we get
@@
It is very important that we put the operator $T_{k}$ on the right side in Proposition 3.4 and not on the left. In fact, the operator $\tilde{S}_{n}=\frac{1}{n}\sum_{k=0}^{n}(1-q)^{k}\tilde{T}_{k}c(\xi)^{k}c(\eta)^{k}$ need not converge to 00 for fixed unit vectors $\xi,\eta\in\mathsf{H}_{U}$ with $\langle\xi,\eta\rangle_{U}=0$ and bounded operators $\tilde{T}_{k}$. See for example Proposition 3.8 below.
We now restrict our attention to a two-dimensional invariant subspace of a given representation $\{U_{t}\}_{t\in\mathbb{R}}$. Let $\mathsf{K}_{\mathbb{R}}=\mathbb{R}^{2}$ be an invariant subspace of $\{U_{t}\}_{t\in\mathbb{R}}$ such that ${U_{t}}_{|_{\mathsf{K}_{\mathbb{R}}}}$ is ergodic i.e. has no fixed nonzero vector. Then, as discussed in Section 2, there is a scalar $\lambda\in(0,1)$ such that
@@
Moreover, the (right) conjugate of $e$ and $\bar{e}$ with respect to the decomposition in $\mathsf{H}_{\mathbb{R}}^{\prime}+i\mathsf{H}_{\mathbb{R}}^{\prime}$ is given as follows:
@@
We briefly sketch the proof from [BMRW, Lemma 3.8] in order to set some notations and record some useful observations. Note that both the sequences have the same norm; indeed we have $W_{r}(\bar{e}^{n}e^{n})=J_{\varphi}W(\bar{e}^{n}e^{n})J_{\varphi}$. We consider the first sequence which by Wick formula is given by
@@
where $q_{k,\ell}^{(n)}=\sum_{J_{1},J_{2}}q^{c(J,J^{c})}$ with the sum running over all possible subsets $J_{1}\subseteq\{1,\ldots,n\}$ of cardinality $k$ and $J_{2}\subseteq\{n+1,\ldots,2n\}$ of cardinality $\ell$ such that $J=J_{1}\cup J_{2}$. Here $c(J,J^{c})$ denotes the number of crossings of the partition $J\cup J^{c}$ of $\{1,2,\ldots,2n\}$ (see [BMRW, Lemma 3.8] for more explanation of the terms $q_{k,\ell}^{(n)}$). Then we have $|q_{k,\ell}^{(n)}|\leq C_{q}^{2}|q|^{(n-k)\ell}$ and
@@
$s_{n}$ $=\frac{1}{n}\sum_{m=1}^{n}\sum_{0\leq k<\ell\leq m}q^{(m-k)\ell}\lambda^{\frac{k-\ell}{2}}+\frac{1}{n}\sum_{m=1}^{n}\sum_{1\leq\ell\leq k\leq m-1}q^{(m-k)\ell}\lambda^{\frac{k-\ell}{2}}$
@@
where $C(q,\lambda,n_{0})$ is a constant independent from $n$, and where we have also used the fact that the sequence $\left\{(m+1)\frac{q^{m}}{2q^{n_{0}}-q^{m}}\right\}_{m\geq 1}$ converges to 00 as $m\to\infty$ and hence its Cesáro sum also converges to 00. On the other hand, the second quantity in (5) is estimated as follows:
@@
as the sequence $\{\frac{q^{k}}{1-q^{k}}\}_{k\geq 1}$ converges to 00 as $k\to\infty$, and so does its Cesáro sum. Thus we have proved that $s_{n}\to 0$ as $n\to\infty$. ∎
@@
By [BMRW, Lemma 3.4], the operator $T$ exists and is both positive and invertible. First check that the series defined in (6) above is norm convergent so that $S$ is a well-defined bounded operator. Indeed if we set $Z_{k}=\frac{d_{\infty}}{d_{k}}(1-q)^{k}\;\lambda^{\frac{k}{2}}c(\bar{e})^{k}Tc(e)^{k}$ for $k\geq 0$ then by using Lemma 2.5 we have
$\sum_{k=0}^{\infty}\|Z_{k}\|$ $\leq\sum_{k=0}^{\infty}\;\frac{d_{\infty}}{d_{k}}(1-q)^{k}\;\lambda^{\frac{k}{2}}\|c(\bar{e})^{k}\|\;\|T\|\;\|c(e)^{k}\|$
@@
where $q_{k,\ell}^{(m)}$ are the scalars as in the proof of Lemma 3.6 which satisfy $|q^{(m)}_{k,\ell}|\leq C_{q}^{2}|q|^{(m-k)\ell}$. Let us understand the terms $c^{*}(\bar{e})^{m-\ell}c(\bar{e})^{m}c(e)^{m}$ in the sum above. Use Corollary 3.2 to write
@@
$\leq\frac{C_{q}^{9}A_{q}}{n^{2}}\sum_{n\geq m>m^{\prime}\geq 1}(1-q)^{2m+2m^{\prime}}\sum_{\ell=0}^{m}\sum_{\ell^{\prime}=0}^{m^{\prime}}|q|^{\ell^{\prime}}\frac{d_{m}d_{m^{\prime}}(1-q)^{-m-m^{\prime}}}{d_{\ell}d_{\ell^{\prime}}(1-q)^{-\ell-\ell^{\prime}}}\lambda^{\frac{m+m^{\prime}-\ell-\ell^{\prime}}{2}}(1-q)^{-m-m^{\prime}-\ell-\ell^{\prime}}\quad(\text{as }q_{m,\ell}^{(m)}\leq C_{q}^{2})$
@@
where $T_{p}=(1-q)^{p}\lambda^{\frac{p}{2}}c^{*}(e)^{p}c(e)^{p}$ for any $p\geq 1$. Let $T=\lim_{p\to\infty}T_{p}$ in norm (which exists by [BMRW, Lemma 3.4] as mentioned above). Also for $0\leq k\leq m$, it is straightforward to verify that $q_{k,0}^{(m)}=\begin{pmatrix}m\\ k\end{pmatrix}_{q}$ (see for example, [BKS, Corollary 2.8]); so that for each $k\geq 0$, we have
@@
In particular, the Cesáro sum of the above sequence has the same limit; hence we get
@@
where we have used the fact that $\|T^{-1}\|\;\|T\|\leq\frac{1}{d_{\infty}}\leq C_{q}$ from [BMRW, Lemma 3.4]. In particular, $d_{\infty}^{-1}T^{-1}S$ is invertible if $C_{q}^{4}\frac{\lambda^{\frac{1}{2}}}{{1-\lambda^{\frac{1}{2}}}}<1$, which is satisfied by the given value of $\lambda$ in the hypothesis. ∎
@@
and since $T^{-1}\leq 2d_{\infty}^{-1}\leq 2C_{q}$ ([BMRW, Lemma 3.4]), we must have
@@
2. Although we have shown that the Cesáro sum of the sequence $\{W(\bar{e}^{m}e^{m})c(\bar{e})^{m}c(e)^{m}\}_{m\geq 1}$ converges, the sequence itself does not converge. For example, this can be checked for $q=0$ where the main obstruction is that the sequence $\{\sum_{\ell=0}^{m}c(\bar{e})^{m}c(e)^{\ell}c^{*}(\bar{e})^{m-\ell}c(\bar{e})^{m}c(e)^{m}\}_{m\geq 1}$ is not convergent. This is one very important reason for us to consider Cesáro sum of such bounded sequence.
We now prove some convergence results involving both the left and right creation/annihilation operators. For this purpose, we introduce the operator $Q$ defined by
@@
We may and we will assume that $\langle\xi,\eta\rangle_{U}\neq 0$, otherwise there is nothing to prove since $r^{*}(\xi)c(\eta)=c(\eta)r^{*}(\xi)$ whenever $\langle\xi,\eta\rangle_{U}=0$. By replacing $\xi$ with $\frac{\xi}{\langle\eta,\xi\rangle_{U}}$ (or $\eta$ with $\frac{\eta}{\langle\xi,\eta\rangle_{U}})$, we will further assume that $\langle\xi,\eta\rangle_{U}=1$.
@@
$=q^{n-1}c(\eta)^{n-1}Q+c(\eta)\left(c(\eta)^{n-1}r^{*}(\xi)+[n-1]_{q}c(\eta)^{n-2}Q\right)$
$=c(\eta)^{n}r^{*}(\xi)+[n]_{q}c(\eta)^{n-1}Q$
as required. Now we prove the result for general $m$ and $n$ such that $n\geq m$. The case $n<m$ follows similarly. So now we fix $n$. The proof proceeds by induction on $m$. For $m=1$, the expression is proved above. By assuming the expression to be true for $m-1$, we calculate
@@
The following proposition can be considered a stronger version of [BMRW, Lemma 3.11] in the sense of norm convergence replacing the weak convergence.
@@
and $P_{\Omega}$ denotes the projection onto the vacuum space $\mathbb{C}\Omega$.
@@
where the sum is divided over $(k,\ell=m)$, $(k,\ell=0)$ and the rest, i.e. $X_{n}$ is the operator
@@
Thus we have divided $R_{n}$ as a sum of three terms, and we will calculate norm limit of each term. To get the limit of the first term, first verify that $Q^{2m}$ converges to $P_{\Omega}$ as $m\to\infty$, and therefore the average $\frac{1}{n}\sum_{m=1}^{n}d_{m}^{2}Q^{2m}$ converges to $d_{\infty}^{2}P_{\Omega}$ in norm. The second term
@@
The sequence $\{m^{2}q^{m}\}_{m\geq 1}$ converges to 00 as $m\to\infty$, and so does its average; hence we conclude that $X_{n}$ converges to 00 in norm as $n\to\infty$. ∎
@@
where $q_{k,\ell}^{(m)}$ are the same scalars as in the proof of Lemma 3.6.
@@ 4. Main result
The quotient $\mathsf{M}^{\omega}:=\prod_{n\geq 1}\mathsf{M}/I_{\omega}$ is a well-defined von Neumann algebra, called ultraproduct of $\mathsf{M}$. A typical element in $\mathsf{M}^{\omega}$ will be denoted as $(x_{n})_{\omega}$ which is the equivalence class of the sequence $(x_{n})$ in the quotient. It is known that $\mathsf{M}$ is a full factor if and only if $\mathsf{M}^{\prime}\cap\mathsf{M}^{\omega}=\mathbb{C}$ [Con1]. The proof of our main result on fullness uses this particular equivalent condition. In what follows, one can also directly employ the equivalent condition of fullness as formulated in [Mar, Corollary B].
Proof of Theorem 1.1. If $\mathsf{H}_{\mathbb{R}}$ is finite-dimensional, then $\mathsf{M}_{q}$ is full by [KSW, Theorem 2.7]. If $\mathsf{H}_{\mathbb{R}}$ has a non-zero weakly mixing part or if the point spectrum $\sigma_{p}(A)$ of eigenvalues of the analytic generator $A$ of $\{U_{t}\}_{t\in\mathbb{R}}$ is infinite and bounded, then $\mathsf{M}_{q}$ is full by [HI, Theorem 6.2]. So assume that $\mathsf{H}_{\mathbb{R}}$ is infinite-dimensional and there is a strictly decreasing sequence of eigenvalues of $A$ converging to 00.
Now let $\omega$ be a fixed non-principal ultrafilter. Let $(X_{k})_{\omega}\in\mathsf{M}_{q}^{\prime}\cap\mathsf{M}_{q}^{\omega}$. By replacing each $X_{k}$ by $X_{k}-\varphi(X_{k})$, we may assume that $\varphi(X_{k})=0$ for all $k\geq 1$. In particular, we have $X_{k}\Omega,X_{k}^{*}\Omega\in\mathcal{F}_{q}(\mathsf{H}_{U})\ominus\mathbb{C}\Omega$ for all $k\geq 1$. We will show that $(X_{k})_{\omega}=0$ in $\mathsf{M}_{q}^{\omega}$ i.e. $\lim_{k\to\omega}\|X_{k}\Omega\|_{q}=0=\lim_{k\to\omega}\|X_{k}^{*}\Omega\|_{q}$.
@@
Assume that $\lambda<(1+C_{q}^{4})^{-2}$ so that the norm limit $S$ of $S_{n}$ exists and is invertible by Proposition 3.8. We observe
@@
where we could exchange the limit because $S_{n}S^{-1}$ converges in norm to identity as $n\to\infty$. Now we use the expressions $XX_{k}-X_{k}X\to 0$-strongly as $k\to\omega$ for all $X\in\mathsf{M}_{q}$, and $X_{k}X^{\prime}-X^{\prime}X_{k}=0$ for all $X^{\prime}\in\mathsf{M}_{q}^{\prime}$ to get
@@
Propositions 3.4 and 3.11 tell us that the operators $R_{n}$ (or $R_{n}^{*}$) when restricted to $\mathcal{F}_{q}(\mathsf{H}_{U})\ominus\mathbb{C}\Omega$ converge to 00 in norm as $n\to\infty$. Since $X_{k}\Omega\in\mathcal{F}_{q}(\mathsf{H}_{U})\ominus\mathbb{C}\Omega$, it follows that
@@
where $L=\sup_{k\geq 1}\|X_{k}\Omega\|^{2}$, and we have used the upper bound for $\|S^{-1}\|$ from Remark 3.9. By our assumption, we have a sequence of such $\lambda$’s which converges to 00; hence it follows that $\lim_{k\to\omega}\|X_{k}\Omega\|_{q}=0$. A similar argument shows that $\lim_{k\to\omega}\|X_{k}^{*}\Omega\|_{q}=0$. This completes the proof.
Acknowledgments. We are grateful to Adam Skalski and Zhiyuan Yang for going through the initial draft, and providing valuable suggestions; particularly, Zhiyuan suggested a shortened proof of Proposition 3.11 by utilizing the operator $Q$. We thank the referee for their careful reading of the manuscript and for several useful comments. This work was partially completed when M.K. was visiting S.W. at Harbin Institute of Technology in Aug-Sep 2023 for a month. M.K. extends appreciation to HIT for their warm hospitality and to S.W. for the invitation and financial support.
M.K. was partially supported by the National Science Center (NCN) grant no. 2020/39/I/ST1/01566. The project is co-financed by the Polish National Agency for Academic Exchange within Polish Returns Program for Mateusz Wasilewski. S.W. was partially supported by the NSF of China (No.12031004, No.12301161) and the Fundamental Research Funds for the Central Universities.